# Reliable Statistical Guarantees for Conformal Predictors with Small Datasets

### v1

@@ Abstract
Surrogate models –including deep neural networks and other machine learning algorithms in supervised learning– are capable of approximating arbitrarily complex, high-dimensional input-output problems in science and engineering, but require a thorough data-agnostic uncertainty quantification analysis before these can be deployed for any safety-critical application. The standard approach for data-agnostic uncertainty quantification is to use conformal prediction (CP), a well-established framework to build uncertainty models with proven statistical guarantees that do not assume any shape for the error distribution of the surrogate model. However, since the classic statistical guarantee offered by CP is given in terms of bounds for the marginal (i.e. ensemble-averaged) coverage, for small calibration set sizes –which are frequent in realistic surrogate modelling that aims to quantify error at different regions–, the potentially strong dispersion of the coverage distribution around its average negatively impacts the reliability of the uncertainty model, often obtaining coverages below the expected value, resulting in a less applicable framework. After providing a gentle presentation of uncertainty quantification for surrogate models for machine learning practitioners, in this paper we bridge the gap by proposing a new statistical guarantee that offers probabilistic information for the coverage of a single conformal predictor. We show that the proposed framework converges to the standard solution offered by CP for large calibration set sizes and, unlike the classic guarantee, still offers reliable information about the coverage of a conformal predictor for small data sizes. We illustrate and validate the methodology in a suite of simple and pedagogical examples, and implement an open access, user-friendly software solution that can be used alongside common conformal prediction libraries to obtain uncertainty models that fulfil the new guarantee.
@@ 1 Introduction
Now, the process of building a conformal predictor from a given UM can conceptually be seen as adding a small correction to the output of the original UM, so that the corrected output achieves some guarantees. Despite their much-deserved success, the statistical guarantee that a conformal predictor provides usually loses relevance when the dataset used to build the uncertainty model is small. Unfortunately, this is often the case when e.g. one would need to estimate the error attached to a given prediction conditioned on a specific region of the input space where such prediction is made. For instance, when predicting stress-related failure probability in aircraft design [lacasa2025towards], one would like to be able to quantify the error of failure probability for the different regions of the aircraft, but often very concrete regions are only sparsely sampled within the (very costly) dataset. This, in practice, poses a strong limitation for conformal predictors that aim to quantify the location-based uncertainty of the surrogate prediction. In this work we also aim to bridge this gap by proposing a modification in the way a given UM is conformalised. Such proposed methodology constitutes the main technical contribution of this work. We will show that the method matches the guarantees offered by standard conformal prediction for large datasets and, unlike standard conformal prediction, our approach is still able to achieve reliable statistical guarantees even with small datasets, thereby guaranteeing the reliability of the UM regardless of the dataset size we have and potentially strengthening its application in various domains of science, engineering, and industry.
The rest of the paper is organized as follows: section 2 gently introduces the basics of surrogate modelling and of uncertainty models, along with some notation and illustrative examples. We also introduce the polysemic notion of coverage as the main metric used to assess the quality and performance of an uncertainty model. section 3 provides a more in-depth tutorial of the nuances of the coverage, coverage distribution, nominal coverage and marginal coverage. After these, in section 4 we describe the standard conformal prediction framework, the current statistical guarantees it offers and flag their limitation for small datasets. The main novel technical contribution of this work is then presented in section 5, where we introduce a new statistical guarantee for conformal predictors that remains reliable in small datasets, and propose a method to build conformal predictors that fulfil this new guarantee. We validate the methodology and illustrate it in a suite of examples, and also implement the method in an open access Python software [repo]. Finally, in section 6 we offer some concluding remarks and discuss possible applications of this work.
@@ 2.1 Surrogate Modelling
A thorough quantification of the prediction errors in unseen data is especially important in safety-critical domains such as aerospace engineering, where the deployment of surrogate models require a rigorous process of certification that goes beyond global error characterisation [lacasa2025towards]. In this context, the surrogate model’s outputs should not be just point predictions, but prediction sets that take quantify the goodness of the approximation –and the uncertainty around it– the surrogate model is doing. These prediction sets (which, in the case of regression, are called prediction intervals) offer a collection of possible outputs (somewhat centered at $\hat{\mathbf{Y}}_{i}$) where we expect the ground truth $\mathbf{Y}_{i}$ to be. Following the previous examples, in the case of the classification of images of animals, if we are presented with the image of a big, grey canine, one prediction set could be $\left\{\text{wolf, dog}\right\}$, and in the case of the pendulum a prediction interval could be $\theta\in\left[-0.1,0.2\right]\text{rad}$, $\dot{\theta}\in\left[0.1,0.15\right]\text{rad/s}$.
@@ 2.2 Uncertainty Models
To summarize, a valid UM needs not only to be informative (i.e. the size of the prediction sets $\mathscr{S}$ need to be small enough so that the UM actually provides non-trivial information of the uncertainty around each prediction of the surrogate model), but also needs to be reliable: the UM must fulfil some kind of statistical guarantee. Traditionally, this guarantee is provided in terms of its marginal coverage, so that it is as close as possible to the desired marginal coverage specified by the user. Validation of this requirement is usually done by direct inspection of the coverage of the UM measured in a test set. Since the guarantee is expressed in terms of marginalised quantities over an ensemble of calibration sets (i.e. the guarantee provides information about the expected coverage and not about the actual coverage a specific UM will get), we advance that the relevance of these statistical guarantees will be challenged when the size of the calibration set sampling is small, as in this case deviations from the average behaviour can be significant.
@@ 3.1 Sample quantile estimators
In the example presented in subsection 2.3 we enumerated the key steps to build a UM, being one of the steps the estimation of a sample quantile. In what follows, we will consider that the sample quantiles are estimated using order statistics, this being one of the simplest and most used approaches within uncertainty quantification and conformal prediction [^2] The choice of the quantile estimator is highly relevant to the properties of the UM coverage. Since the conformal prediction framework assumes that the quantiles are computed with order statistics, we will use this method.. With this in mind, given a generic set (a sample) of $N$ points $\left\{s_{1},s_{2},...,s_{N}\right\}$, the sample quantile at level $q$ is defined as
@@ 4.2 What happens when $N_{\text{cal}}$ is small?
The statistical guarantee of a conformal predictor (given by Equation 13) is given in terms of bounds for the marginal coverage (i.e., the expected value of the coverage). However, as illustrated in Figure 3, for sufficiently small calibration set sizes, the bounds provided by the guarantee are not representative of the actual coverage a conformal predictor will get: the spread of the coverage distribution for small values of $N_{\text{cal}}$ can cause the obtained coverage to be much smaller than the lower bound of the marginal coverage in a considerable proportion of cases. This is specially detrimental in cases where reliability is key (as it happens in safety-critical applications), and where data is limited or costly to acquire, rendering small calibration set sizes.
Importantly, observe that even if the original dataset is large, it is often the case that we need to conformalise the uncertainty model for different regions of the input space $S_{X}$. This is called group-balanced conformal prediction, and it seeks to obtain a conformal predictor that fulfils Equation 13 for a set of groups of $\mathbf{X}$. For example, if $\mathbf{X}$ is continuous, we would bin $\mathbf{X}$ and, by performing group-balanced conformal prediction, we would fulfil Equation 13 in every bin independently. In contrast, in conformal prediction as described in this section, we would fulfil Equation 13 in $S_{X}$ (globally), but if we evaluate the coverage in a given region of $S_{X}$, the coverage might not fulfil the guarantee. This is sensible to do when prediction errors are heterogeneously distributed throughout the input space, and is often the case when building surrogate models of physical models that for instance show strong gradients over their spatial domain. For all these reasons, being able to build UMs with reliable guarantees even for small calibration sets is certainly needed. In section 5 we will propose a modification to conformal prediction to precisely achieve a statistical guarantee that is still reliable when the dataset size is small, thus producing conformal predictors that are reliable regardless of the calibration set size they are built on.
@@ 4.3 A simple example
In figures 4(c) and 4(d) we show the results for the same procedure repeated using a smaller calibration set of size $N_{\text{cal}}=100$. We can clearly see that the dispersion of the coverage distribution has considerably increased, thus making the bounds for the marginal coverage provided by the conformal prediction guarantee non-reliable indicators of the actual coverage a conformalised UM will achieve.
@@
5 Towards reliable statistical guarantees with small datasets
As seen in the previous sections, the statistical guarantee provided by conformal predictors becomes unreliable when the calibration set size is small. Now, we aim to find a new guarantee that (i) matches the standard one provided by conformal prediction for large datasets, and (ii) unlike the standard approach, still provides useful information about the coverage even with small datasets. To that aim, we define a minimum acceptable coverage $C^{\text{min}}$ and build conformal predictors whose coverage $C$ is at least $C^{\text{min}}$ with a confidence of (at least) a minimum level specified by the user. Formally, we are looking for UMs that satisfy:
@@ 5 Towards reliable statistical guarantees with small datasets
The key difference with respect to the classic conformal prediction guarantee (Equation 13) is that in the classic guarantee we specify bounds for the expected value of the coverage, i.e. the marginal coverage. But we have already pointed out that, with small calibration sets, even if the expected coverage fulfils Equation 13, the spread of the coverage distribution can cause the coverage obtained from a single conformal predictor to deviate considerably from these bounds, rendering the information this guarantee provides unreliable with small datasets. On the other hand, with the new guarantee, we are establishing the minimum coverage we will achieve (with a given confidence) independently of the calibration set size, thus putting more emphasis on the reliability of the UM.
However, as already noted, when the calibration set is small, even if the expected coverage still fulfils Equation 13, the large dispersion of the coverage distribution can cause the coverage of a single conformal predictor to deviate substantially from these bounds, making the guarantee effectively irrelevant in small datasets. On the other hand, with the new guarantee, we are establishing the minimum coverage we will achieve (with a given confidence) independently of the calibration set size, thus putting more emphasis on the reliability of the UM.
@@ 5.1 An illustrative example using a simple UM
To illustrate the differences between a classical conformal with the guarantee Equation 13 and one using our proposed guarantee Equation 14, we will continue to use the example of subsection 4.3. In Figure 5 we compare the coverage distribution of a conformal predictor built using the classic guarantee (blue) compared with the new guarantee (orange), for two calibration set sizes, $N_{\text{cal}}=1000$ (top plots of the figure) and $N_{\text{cal}}=100$ (bottom plots of the figure). From these figures, it is clear that the new guarantee “shifts” the coverage distribution towards higher coverage values, enough so that we can ensure that any conformal predictor will have a coverage above $C^{\text{min}}$ with at least a confidence of $1-\alpha$. This contrasts greatly with the coverage distribution provided by the classic guarantee, where a substantial proportion of the distribution is in a range of coverages lower than the desired $C^{\text{nom}}$, which can lead to conformal predictors with a coverage considerably lower than $C^{\text{nom}}$ when using small calibration sets. This is precisely what happens in the $N_{\text{cal}}=100$ case. Observe in 5(d) that the spread of the coverage distribution is quite large for the classic guarantee. This makes the conformal predictor non reliable: using Equation 13 we know that the marginal coverage $\mathbb{E}(C)$ will be bounded in $\mathbb{E}(C)\in\left[C^{\text{nom}},C^{\text{nom}}+1/(N_{\text{cal}}+1)\right]\approx\left[0.9,0.9099\right]$, but a considerable proportion of conformal predictors will end up having an actual coverage lower than 0.9. Specifically, around 46% will have an actual coverage lower than 0.9, and around 10% will have an actual coverage lower than 0.86. This loss of reliability of the classical guarantee is more acute when the calibration set is smaller, something that does not affect our newly proposed guarantee.
@@ 5.2 On the informativeness of the new guarantee
As seen in the previous section, by enforcing the new guarantee given by Equation 14 we achieve uncertainty models with reliable statistical guarantees even with small samples, where the classical conformalisation fails. However, these newly conformalised UMs are also more conservative than those obtained from classic conformal prediction, i.e. prediction intervals might be larger, potentially challenging the UM’s informative character. While there is no way of knowing a priori if this new type of conformal predictor will be sufficiently informative for a given minimum coverage and confidence (this must be evaluated a posteriori, e.g. by looking at the width of the prediction intervals in a regression problem), this potential loss of informativeness can be easily removed by lowering either the minimum acceptable coverage $C^{\text{min}}$ or the confidence $1-\alpha$. In Figure 6 we present the coverage distributions of two alternatives to the UMs built so far with the new guarantee: one with a lower $C^{\text{min}}$, and another one with a lower confidence. Both options still fulfil Equation 14, so they satisfy a reliable guarantee with small calibration sets, and offer slightly less conservative predictions than the UM presented in figures 5(c) and 5(d). Evidently, these relaxations in either $C^{\text{min}}$ or confidence must be carefully assessed in each particular use case.
@@ 6 Discussion
In this work, we have firstly provided a gentle introduction –almost in tutorial form– to the relationship between surrogate models and their uncertainty models, as well as on the conformal prediction framework and how to conformalise uncertainty models. We have highlighted that standard conformal prediction yields non-reliable guarantees when the calibration set sizes are small, and have proposed a new set of guarantees to conformalise uncertainty models that hold independent of the calibration set size and therefore are still reliable in those cases. These novel statistical guarantees –alongside a method to build conformal predictors with these guarantees– represent the main technical contribution of this work.
Concretely, the proposed guarantee ensures that a given conformal predictor will achieve a coverage of at least $C^{\text{min}}$ with at least a confidence of $1-\alpha$. This guarantee is based on analytical expressions so, provided that the hypotheses on which conformal prediction is based are fulfiled, the validity of the information the guarantee provides is ensured when adequately conformalising any arbitrary uncertainty model. The new guarantee renders very similar results to the classic guarantee for large calibration sample sizes and, unlike the classic guarantee, it is also reliable with small sample sizes. This reliability comes at a risk of producing too conservative conformal predictors, but this can be managed by slightly reducing the requirements in terms of minimum acceptable coverage and/or confidence. Conformal predictors built using our new guarantee will place special emphasis on their reliability, making them particularly useful and well-suited to build UMs when data is scarce and/or when the reliability of the UM is essential: these two ingredients are present in safety-critical applications [CoDANN] and in surrogate models of physical processes. In particular, our proposed conformalisation gains much relevance within group-balanced conformal prediction, i.e. when we aim at creating a UM that yields different prediction intervals for different regions of the parameter space. In this case, our method enables us to conformalise such region-dependent UM and obtain a suite of independent conformal predictors conditioned to different regions, where all of them remain reliable even if the region-dependent calibration sets are small.

### v2

@@ Abstract
Surrogate models –including deep neural networks and other machine learning algorithms in supervised learning– are capable of approximating arbitrarily complex, high-dimensional input-output problems in science and engineering, but require a thorough data-agnostic uncertainty quantification analysis before these can be deployed for any safety-critical application. The standard approach for data-agnostic uncertainty quantification is to use conformal prediction (CP), a well-established framework to build uncertainty models with proven statistical guarantees that do not assume any shape for the error distribution of the surrogate model. However, since the classic statistical guarantee offered by CP is given in terms of bounds for the marginal (i.e. ensemble-averaged) coverage, for small calibration set sizes –which are frequent in realistic surrogate modelling whose prediction error needs to be quantified at different regions–, the potentially strong dispersion of the coverage distribution around its average negatively impacts the relevance of the uncertainty model’s statistical guarantee, often obtaining coverages below the expected value, resulting in a less applicable framework. After providing a gentle presentation of uncertainty quantification for surrogate models for machine learning practitioners, in this paper we bridge the gap by proposing a new statistical guarantee that offers probabilistic information for the coverage of a single conformal predictor. We show that the proposed framework converges to the standard solution offered by CP for large calibration set sizes and, unlike the classic guarantee, still offers relevant information about the coverage of a conformal predictor for small data sizes. We illustrate and validate the methodology in a suite of simple and pedagogical examples, and implement an open access, user-friendly software solution that can be used alongside common conformal prediction libraries to obtain uncertainty models that fulfil the new guarantee.
@@ 1 Introduction
Now, the process of building a conformal predictor from a given UM can conceptually be seen as adding a small correction to the output of the original UM, so that the corrected output achieves some guarantees. Despite their much-deserved success, the statistical guarantee that a conformal predictor provides usually loses relevance when the dataset used to build the uncertainty model is small. Unfortunately, this is often the case when e.g. one would need to estimate the error attached to a given prediction conditioned on a specific region of the input space where such prediction is made. For instance, when predicting stress-related failure probability in aircraft design [lacasa2025towards], one would like to be able to quantify the error associated to a certain prediction of failure probability for each specific region of the aircraft. Unfortunately, often very concrete regions are only sparsely sampled within the (very costly) dataset. This, in practice, poses a strong limitation for conformal predictors that aim to quantify the location-based uncertainty of the surrogate prediction. In this work we also aim to bridge this gap by proposing a modification in the way a given UM is conformalised. Such proposed methodology constitutes the main technical contribution of this work. We will show that the method matches the guarantees offered by standard conformal prediction for large datasets and, unlike standard conformal prediction, our approach is still able to achieve relevant statistical guarantees even with small datasets, thereby guaranteeing the relevance of the UM regardless of the dataset size we have and potentially strengthening its application in various domains of science, engineering, and industry.
The rest of the paper is organized as follows: section 2 gently introduces the basics of surrogate modelling and of uncertainty models, along with some notation and illustrative examples. We also introduce the polysemic notion of coverage as the main metric used to assess the quality and performance of an uncertainty model. section 3 provides a more in-depth tutorial of the nuances of the coverage, coverage distribution, nominal coverage and marginal coverage. After these, in section 4 we describe the standard conformal prediction framework, the current statistical guarantees it offers and flag their limitation for small datasets. The main novel technical contribution of this work is then presented in section 5, where we introduce a new statistical guarantee for conformal predictors that remains relevant in small datasets, and propose a method to build conformal predictors that fulfil this new guarantee. We validate the methodology and illustrate it in a suite of examples, and also implement the method in an open access Python software [repo]. Finally, in section 6 we offer some concluding remarks and discuss possible applications of this work.
@@ 2.1 Surrogate Modelling
A thorough quantification of the prediction errors in unseen data is especially important in safety-critical domains such as aerospace engineering, where the deployment of surrogate models require a rigorous process of certification that goes beyond global error characterisation [lacasa2025towards]. In this context, the surrogate model’s outputs should not be just point predictions, but prediction sets that quantify the goodness of the approximation –and the uncertainty around it– the surrogate model is doing. These prediction sets (which, in the case of regression, are called prediction intervals) offer a collection of possible outputs (somewhat centered at $\hat{\mathbf{Y}}_{i}$) where we expect the ground truth $\mathbf{Y}_{i}$ to be. Following the previous examples, in the case of the classification of images of animals, if we are presented with the image of a big, grey canine, one prediction set could be $\left\{\text{wolf, dog}\right\}$, and in the case of the pendulum a prediction interval could be $\theta\in\left[-0.1,0.2\right]\text{rad}$, $\dot{\theta}\in\left[0.1,0.15\right]\text{rad/s}$.
@@ 2.2 Uncertainty Models
To summarize, a valid UM needs not only to be informative, i.e. the size of the prediction sets $\mathscr{S}$ need to be small enough so that the UM actually provides non-trivial information of the uncertainty around each prediction of the surrogate model (note sometimes informativeness is also called efficiency), but also needs to be reliable: the UM must fulfil some kind of statistical guarantee. Traditionally, this guarantee is provided in terms of its marginal coverage, so that it is as close as possible to the desired marginal coverage specified by the user. Validation of this requirement is usually done by direct inspection of the coverage of the UM measured in a test set. Since the guarantee is expressed in terms of marginalised quantities over an ensemble of calibration sets (i.e. the guarantee provides information about the expected coverage and not about the actual coverage a specific UM will get), we advance that the relevance of these statistical guarantees will be challenged when the size of the calibration set sampling is small, as in this case deviations from the average behaviour can be significant.
@@ 3.1 Sample quantile estimators
In the example presented in subsection 2.3 we enumerated the key steps to build a UM, being one of the steps the estimation of a sample quantile. In what follows, we will consider that the sample quantiles are estimated using order statistics, this being one of the simplest and most used approaches within uncertainty quantification and conformal prediction [^2] The properties of the UM coverage are highly dependant on the choice of the quantile estimator. Since the conformal prediction framework assumes that the quantiles are computed with order statistics, we will use this method.. With this in mind, given a generic set (a sample) of $N$ points $\left\{s_{1},s_{2},...,s_{N}\right\}$, the sample quantile at level $q$ is defined as
@@ 4.2 What happens when $N_{\text{cal}}$ is small?
The statistical guarantee of a conformal predictor (given by Equation 13) is given in terms of bounds for the marginal coverage (i.e., the expected value of the coverage). However, as illustrated in Figure 3, for sufficiently small calibration set sizes, the bounds provided by the guarantee are not representative of the actual coverage a conformal predictor will get: the spread of the coverage distribution for small values of $N_{\text{cal}}$ can cause the obtained coverage to be much smaller than the lower bound of the marginal coverage in a considerable proportion of cases. This is specially detrimental in cases where achieving satisfactory coverages for a single UM is key (as it happens in safety-critical applications), and where data is limited or costly to acquire, rendering small calibration set sizes.
Importantly, observe that even if the original dataset is large, it is often the case that we need to conformalise the uncertainty model for different regions of the input space $S_{X}$. This is called group-balanced conformal prediction, and it seeks to obtain a conformal predictor that fulfils Equation 13 for a set of groups of $\mathbf{X}$. For example, if $\mathbf{X}$ is continuous, we would bin $\mathbf{X}$ and, by performing group-balanced conformal prediction, we would fulfil Equation 13 in every bin independently. In contrast, in conformal prediction as described in this section, we would fulfil Equation 13 in $S_{X}$ (globally), but if we evaluate the coverage in a given region of $S_{X}$, the coverage might not fulfil the guarantee. This is sensible to do when prediction errors are heterogeneously distributed throughout the input space, and is often the case when building surrogate models of physical models that for instance show strong gradients over their spatial domain. For all these reasons, being able to build UMs with relevant guarantees even for small calibration sets is certainly needed. In section 5 we will propose a modification to conformal prediction to precisely achieve a statistical guarantee that is still relevant when the dataset size is small, thus producing conformal predictors that provide useful prediction intervals regardless of the calibration set size they are built on.
@@ 4.3 A simple example
In figures 4(c) and 4(d) we show the results for the same procedure repeated using a smaller calibration set of size $N_{\text{cal}}=100$. We can clearly see that the dispersion of the coverage distribution has considerably increased, thus making the bounds for the marginal coverage provided by the conformal prediction guarantee non-relevant as indicators of the actual coverage a conformalised UM will achieve.
@@
5 Towards relevant statistical guarantees with small datasets
As seen in the previous sections, the statistical guarantee provided by conformal predictors becomes non-relevant when the calibration set size is small. Now, we aim to find a new guarantee that (i) matches the standard one provided by conformal prediction for large datasets, and (ii) unlike the standard approach, still provides useful information about the coverage even with small datasets. To that aim, we define a minimum acceptable coverage $C^{\text{min}}$ and build conformal predictors whose coverage $C$ is at least $C^{\text{min}}$ with a confidence of (at least) a minimum level specified by the user. Formally, we are looking for UMs that satisfy:
@@ 5 Towards relevant statistical guarantees with small datasets
The key difference with respect to the classic conformal prediction guarantee (Equation 13) is that in the classic guarantee we specify bounds for the expected value of the coverage, i.e. the marginal coverage. However, as already noted, when the calibration set is small, even if the expected coverage still fulfils Equation 13, the large dispersion of the coverage distribution can cause the coverage of a single conformal predictor to deviate substantially from these bounds, making the classic guarantee effectively irrelevant in small datasets. On the other hand, with the new guarantee, we are establishing the minimum coverage we will achieve (with a given confidence) independently of the calibration set size, thus putting more emphasis on the relevance of the statistical guarantee of the UM.
@@ 5.1 An illustrative example using a simple UM
To illustrate the differences between a classical conformal with the guarantee Equation 13 and one using our proposed guarantee Equation 14, we will continue to use the example of subsection 4.3. In Figure 5 we compare the coverage distribution of a conformal predictor built using the classic guarantee (blue) compared with the new guarantee (orange), for two calibration set sizes, $N_{\text{cal}}=1000$ (top plots of the figure) and $N_{\text{cal}}=100$ (bottom plots of the figure). From these figures, it is clear that the new guarantee “shifts” the coverage distribution towards higher coverage values, enough so that we can ensure that any conformal predictor will have a coverage above $C^{\text{min}}$ with at least a confidence of $1-\alpha$. This contrasts greatly with the coverage distribution provided by the classic guarantee, where a substantial proportion of the distribution is in a range of coverages lower than the desired $C^{\text{nom}}$, which can lead to conformal predictors with a coverage considerably lower than $C^{\text{nom}}$ when using small calibration sets. This is precisely what happens in the $N_{\text{cal}}=100$ case. Observe in 5(d) that the spread of the coverage distribution is quite large for the classic guarantee. This makes the conformal predictor’s statistical guarantee non-relevant: using Equation 13 we know that the marginal coverage $\mathbb{E}(C)$ will be bounded in $\mathbb{E}(C)\in\left[C^{\text{nom}},C^{\text{nom}}+1/(N_{\text{cal}}+1)\right]\approx\left[0.9,0.9099\right]$, but a considerable proportion of conformal predictors will end up having an actual coverage lower than 0.9. Specifically, around 46% will have an actual coverage lower than 0.9, and around 10% will have an actual coverage lower than 0.86. This loss of relevance of the classical guarantee is more acute when the calibration set is smaller, something that does not affect our newly proposed guarantee.
@@ 5.2 On the informativeness of the new guarantee
As seen in the previous section, by enforcing the new guarantee given by Equation 14 we achieve uncertainty models with relevant statistical guarantees even with small samples, where the classical conformalisation fails. However, these newly conformalised UMs are also more conservative than those obtained from classic conformal prediction, i.e. prediction intervals might be larger, potentially challenging the UM’s informative character. While there is no way of knowing a priori if this new type of conformal predictor will be sufficiently informative for a given minimum coverage and confidence (this must be evaluated a posteriori, e.g. by looking at the width of the prediction intervals in a regression problem), this potential loss of informativeness can be easily removed by lowering either the minimum acceptable coverage $C^{\text{min}}$ or the confidence $1-\alpha$. In Figure 6 we present the coverage distributions of two alternatives to the UMs built so far with the new guarantee: one with a lower $C^{\text{min}}$, and another one with a lower confidence. Both options still fulfil Equation 14, so they satisfy a relevant guarantee with small calibration sets, and offer slightly less conservative predictions than the UM presented in figures 5(c) and 5(d). Evidently, these relaxations in either $C^{\text{min}}$ or confidence must be carefully assessed in each particular use case.
@@ 6 Discussion
In this work, we have firstly provided a gentle introduction –almost in tutorial form– to the relationship between surrogate models and their uncertainty models, as well as on the conformal prediction framework and how to conformalise uncertainty models. We have highlighted that standard conformal prediction yields non-relevant guarantees when the calibration set sizes are small, and have proposed a new set of guarantees to conformalise uncertainty models that are relevant independently of the calibration set size and therefore are still useful in those cases. These novel statistical guarantees –alongside a method to build conformal predictors with these guarantees– represent the main technical contribution of this work.
Concretely, the proposed guarantee ensures that a given conformal predictor will achieve a coverage of at least $C^{\text{min}}$ with at least a confidence of $1-\alpha$. This guarantee is based on analytical expressions so, provided that the hypotheses on which conformal prediction is based are fulfilled, the validity of the information the guarantee provides is ensured when adequately conformalising any arbitrary uncertainty model. The new guarantee renders very similar results to the classic guarantee for large calibration sample sizes and, unlike the classic guarantee, it is also relevant with small sample sizes. This relevance comes at a risk of producing too conservative conformal predictors, but this can be managed by slightly reducing the requirements in terms of minimum acceptable coverage and/or confidence. Conformal predictors built using our new guarantee will place special emphasis on the relevance of the guarantees they provide, making them particularly useful and well-suited to build UMs when data is scarce and/or when achieving satisfactory coverages for a single UM is essential: these two ingredients are present in safety-critical applications [CoDANN] and in surrogate models of physical processes. In particular, our proposed conformalisation gains much relevance within group-balanced conformal prediction, i.e. when we aim at creating a UM that yields different prediction intervals for different regions of the parameter space. In this case, our method enables us to conformalise such region-dependent UM and obtain a suite of independent conformal predictors conditioned to different regions, where all of them remain useful even if the region-dependent calibration sets are small.